Kirgan, Inc., manufactures a product with the following costs: Per Year Per Unit $26.40 $15.40 $ 3.60 Direct materials DirectMagney, Inc., uses the absorption costing approach to cost-plus pricing described in the text to set prices for its products.
Answer #1

1.

Cost for 96000 units
Direct Material $     2,534,400 =96000*26.4
Direct Labor $     1,478,400 =96000*15.4
Variable Manufacturing Overhead $        345,600 =96000*3.6
Fixed Manufacturing Overhead $     1,545,600
Variable Selling and administrative expenses $        336,000 =96000*3.5
Fixed Selling and administrative expenses $     1,514,700
Total Cost $     7,754,700
Target Profit $           59,200 =370000*16%
Sales Revenue $     7,813,900
Units sold 96000
Selling price per unit $             81.40

Answer is b. $81.40

2.

Product Cost $     1,946,800
Selling and administrative expenses $        829,600
Target Profit $           51,300
Sales Revenue $     2,827,700
Units sold 31000
Selling price per unit $             91.22

Answer is a. $91.22